A monopoly firm faces a demand curve given by the following equation: P = $500 − 10Q, where Q equals quantity sold per day. Its marginal cost curve is MC = $100 per day. Assume that the firm faces no fixed cost. You may wish to arrive at the answers mathematically, or by using a graph (the graph is not required to be presented), either way, please provide a brief description of how you arrived at your results.
a) How much will the firm produce?
1
Expert's answer
2021-09-28T18:07:32-0400

MR = MC, so:

"MR = TR'(Q) = 500 - 20Q, MC = 100,"

"500 - 20Q = 100,"

"Q = 20 units"

= 20units